### v1

# Information Gain-Guided Causal Intervention for Autonomous Debiasing Large Language Models

Abstract
Despite significant progress, recent studies indicate that current large language models (LLMs) may still capture dataset biases and utilize them during inference, leading to the poor generalizability of LLMs. However, due to the diversity of dataset biases and the insufficient nature of bias suppression based on in-context learning, the effectiveness of previous prior knowledge-based debiasing methods and in-context learning based automatic debiasing methods is limited. To address these challenges, we explore the combination of causal mechanisms with information theory and propose an information gain-guided causal intervention debiasing (IGCIDB) framework. This framework first utilizes an information gain-guided causal intervention method to automatically and autonomously balance the distribution of instruction-tuning dataset. Subsequently, it employs a standard supervised fine-tuning process to train LLMs on the debiased dataset. Experimental results show that IGCIDB can effectively debias LLM to improve its generalizability across different tasks.
Information Gain-Guided Causal Intervention for Autonomous Debiasing Large Language Models
Zhouhao Sun1, Xiao Ding1thanks: Corresponding Author, Li Du2, Yunpeng Xu1, Yixuan Ma1, Yang Zhao1, Bing Qin1, Ting Liu1 1Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China 2Beijing Academy of Artificial Intelligence, Beijing, China {zhsun, xding, ypxu, yxma, yzhao, bqin, tliu}@ir.hit.edu.cn duli@baai.ac.cn
1 Introduction
Large language models (LLMs), through the pre-training and supervised fine-tuning process, have demonstrated remarkable ability to follow human instructions to solve various tasks (Achiam et al., 2023), demonstrating immense potential in real-world applications.
Figure 1: The examples in dataset before and after data debiasing. By intervening example 2, answer with less-known individual are introduced (MORGAN Moses) so that LLMs cannot learn to make predictions solely relying on an individual’s popularity.
Although achieving promising performance, LLMs would also inevitably learn dataset bias during the supervised fine-tuning process, such as negation bias and popularity bias Schick et al. (2021); Mukherjee et al. (2021); Navigli et al. (2023); Klimashevskaia et al. (2024). These dataset biases would lead to poor generalizability of LLMs (Yang et al., 2024; Cheng and Amiri, 2024; Yang et al., 2025). As shown in Figure 1, the original dataset only contains samples where the answers are well-known figures, and it lacks samples featuring less-known individuals as answers. Consequently, LLMs trained on such instruction-tuning corpus would also learn to utilize the biased feature ’popularity of the individuals’ for making predictions. However, when less-known people are generally the correct answer in some datasets, the performance of the LLM will significantly decline, indicating that dataset biases would reduce the generalizability of LLMs.
This problem highlights the necessity of debiasing LLMs. While contemporary debiasing methods are relatively effective, most of them require prior knowledge of dataset biases (Oba et al., 2023; Liu et al., 2023; Cheng and Amiri, 2024). However, dataset biases can vary significantly across different datasets and tasks (Poliak et al., 2018; Schuster et al., 2019; Schick et al., 2021), making it impractical to identify them one by one manually. Furthermore, a vast amount of biases still remain unrecognized in different datasets (Nie et al., 2020) and new biases are constantly being discovered.
Hence, there is an urgent demand for methods to automatically debias LLMs without prior knowledge of dataset biases. However, previous automatic debiasing approaches (Sun et al., 2024) for LLMs are based on in-context learning (ICL) which can only suppress bias in LLMs instead of completely eliminating it in theory, as they depend on the instruction to mitigate bias without altering the model’s intrinsic parameters.
To address these issues, we start from the essential cause to debias the instruction-tuning dataset and propose a Information Gain-guided Causal Intervention DeBiasing Framework (IGCIDB). To eliminate biased features within the instruction-tuning dataset, it is essential to ensure that these features do not provide any additional information to predict the answers, i.e., the information gain (Csiszár and Körner, 2015) of these features for predicting the answers needs to be 0. Therefore, IGCIDB initially employs the CAL Sun et al. (2024) method to automatically detect biased features within the datasets. Subsequently, the information gain is utilized to check the relationships between biased features and answers. Following this, IGCIDB utilize a causal intervention-based data rewriting method to intervene in the biased features. This process can modify the joint distribution between biased features and answers to eliminate their correlations, effectively reducing the information gain. Finally, LLMs are fine-tuned on the debiased dataset utilizing the supervised fine-tuning method.
As illustrated in Figure 1, the original dataset is biased because it exclusively contains data whose answer is a well-known individual, making LLMs trained on such dataset rely sorely on the individual’s popularity when answering similar questions. By utilizing the information gain-guided causal intervention method to revise example 2, the proportion of well-known individual (James) and less-known individual (MORGAN Moses) in the answer becomes the same. Consequently, the biased feature individual’s popularity does not provide any information gain to predict the answers which indicates the success of the data debiasing process.
Experimental results show that our approach can improve the generalizability of LLMs on transfer and challenge datasets for different tasks. We also note that dataset debiased by a certain LLM can also be utilized for another LLM to mitigate bias.
2 Biased Features within the Instruction-tuning Datasets
Text records and reflects human thoughts. However, due to inherent prejudice or preference of human beings, as well as potential annotation artifacts, various biases such as gender and popularity biases still broadly exist in various datasets Steen and Markert (2024); Dai et al. (2024).
Figure 2: (a) Three steps of Information Gain-Guided Causal Intervention Debiasing Framework. (b) Illustration of the Information Gain-guided Data Debiasing Goal, i.e., the biased feature cannot contribute to correctly predicting the answers. (c) Illustration of the Causal Intervention-based Data Rewriting process.
In instruction-tuning datasets, each piece of data consists of a question (i.e., instruction) and an answer. And each instruction-tuning dataset contains three types of features: The first category of features provides no assistance in accurately answering the questions within the dataset and is referred to as the non-predictive feature. The remaining features are collectively termed predictive features. Among these predictive features, causal features directly determine the answer and exhibit a causal relationship with it. The other predictive features, which do not have a direct causal relation to the answer, can be defined as predictive but non-causal features. For instance, in many datasets, the presence of negation word such as ‘not’ often appears when the sentiment of the sentence is negative rather than positive (Mukherjee et al., 2021). However, the mere presence of the word ‘not’ does not necessarily indicate that the sentiment is negative. In this case, the presence of the word "not" serves as a predictive but non-causal feature. In the debiasing scenario, the predictive but non-causal feature are also referred to as biased features.
Among these categories of features, non-predictive features are generally not learned and utilized by powerful LLMs to answer questions, as they do not contribute to providing correct answers. Predictive features, which are useful for accurately answering questions within the dataset, are typically learned by LLMs. However, biased features lack a causal relationship with the answers. Considering that data distribution varies across different datasets, the predictive power of these biased features for answers may change. For example, if the sentiment of sentences in a dataset is generally positive when the negation word ‘not’ appears, the performance of LLMs have learned to predict negative sentiment based on the negation word may decline. Consequently, compared to learning causal features, the out-of-distribution generalization ability of LLMs may be decreased by learning and utilizing biased features.
3 Method
As illustrated in Figure 2 (a), IGCIDB contains three main steps: (i) Fine-tune the pre-trained LLM on original biased instruction-tuning dataset; (ii) Biased feature identification and information gain-guided causal intervention for data debiasing. (iii) Fine-tune the pre-trained LLM on unbiased instruction-tuning dataset. Since (i) and (iii) are standard supervised fine-tuning processes, our focus will be on providing a detailed explanation on process (ii).
3.1 Biased Feature Identification
For brevity, we denote the LLM fine-tuned on the original biased instruction-tuning datasets as the biased LLM. To automatically identify explainable biased features, we utilize the causal-guided active learning (CAL) method proposed by Sun et al. (2024). This method is applied independently to each task within the original instruction-tuning dataset. Through this approach, we are able to identify a series of biased features using the biased LLM. Though there are certain limitations, such as being unable to deal with open-ended tasks without standard answers (e.g. writing essays), it is currently the only automated method for identifying interpretable biased features.
Specifically, for discriminative tasks, we directly apply the original CAL method to automatically identify biased features. For other tasks with standard answers such as the QA task, we apply the CAL method to data with consistent question objectives [^1] The consistency of question objectives also means that the answer space of these data to be the same for each category of data to identify biased features, since the CAL method requires uniformity in the question’s goal.
3.2 Information Gain-guided Data Debiasing Goal
To debias the instruction-tuning dataset, we need to ensure that the identified biased features do not provide any additional information to predict the answers. That is, we need to make the information gain (Csiszár and Körner, 2015) of these features for predicting the answers to be 0 under this dataset. Specifically, for any biased feature $B$:
$IG(Y,B)=0.$ (1)
where $IG$ is the information gain, and $Y$ is the answer. And it can be satisfied if and only if:
$P(Y|B)=P(Y).$ (2)
The specific derivation process for Eq. 2 can be found in the Appendix A. Since there is no prior knowledge about $Y$, the probability distribution of $P(Y)$ should be uniform to avoid introducing label bias. Consequently, for any two values $y_{i}$ and $y_{j}$ of $Y$, we have
$P(y_{i}|B)=P(y_{j}|B).$ (3)
As we aims at debiasing the dataset, frequency is employed to replace probability. Consequently, given the dataset, for any value $b_{k}$ of $B$, we have
$N(y_{i},b_{k})=N(y_{j},b_{k}).$ (4)
where $N(y_{i},b_{k})$ denotes the number of instances whose answer and biased feature are $y_{i}$ and $b_{k}$, respectively. The information gain-guided data debiasing goal is shown in Figure 2 (b). Note that even if the number $N$ exhibits slight disparities, LLMs are unlikely to learn these subtle correlations. Therefore, we relax the strict equality in Eq. 4 to an approximate equality. To the best of our knowledge, we are the first to establish strict criteria that an unbiased instruction-tuning dataset must satisfy.
For scenarios where the value space of the biased feature $B$ and answer $Y$ can be easily enumerated, such as the negation bias [^2] A negation word either exists in the sentence or not. in the sentiment analysis task, this data debiasing goal can directly guide the subsequent causal intervention process. However, for scenarios where the value space of the biased feature $B$ or the answer $Y$ is difficult to enumerate, the data debiasing goal should be adjusted. For example, if the biased feature is the individual’s popularity in the given context and the question is who caused a certain event according to the given context, the value space of the biased feature will be the popularity of people appearing in the context, and the value space of the answer will be the people appearing in the context. Since the people appearing in each context are different, it is difficult to enumerate them one by one. Additionally, it does not matter who the people are appearing in the context, but the popularity of the people matters. As a result, the value space of the biased feature can be changed to the ratio of popular people and unpopular people appearing in the context, and the value space of the answer can be changed to the popularity of the answer character. Other data debiasing goals that need to be adjusted and the adjustment process can be found in the Appendix B.
3.3 Causal Intervention-based Data Rewriting
The process of causal intervention (Pearl, 2009) involves modifying the original data distribution to produce a distribution that meets our expectations, which enables the realization of the information gain-guided data debiasing goal. Specifically, we first analyze the original distribution of the dataset and determine the intervention direction (e.g., the arrow in Figure 2 (c)). Subsequently, a causal intervention procedure is applied to a proportion of data by executing the operation $do(B=b_{k})$, thereby modifying the data distribution to meet the data debiasing goal. This causal intervention process intervenes in the biased features, without changing the other semantics of the samples [^3] For semantic that tightly tied to biased features, we also need to modify the semantics of the sample. For example, the individuals are tied to popularity, so we need to modify the figure when modifying popularity..
For instance, as shown in Figure 2 (c), the label’s frequency of the dataset is imbalanced when the negation word exists, suggesting that the existence of the negation word contributes to correctly predicting the answer of the dataset. In this example, $B$ represents the presence or absence of a negation word in a sentence, and we denote $b_{k}$ as the presence of the negation word. If we apply the operation $do(B=b_{k})$ for a fixed number of samples that do not contain negation words and have a positive sentiment, they will be transformed into samples that include negation words while maintaining the same positive sentiment [^4] For example, if the origin sample is ‘you will find yourself remembering this refreshing visit to a sunshine state’, the operation $do(B=Yes)$ can be achieved by replacing the word ‘remembering’ with ‘cannot forget’., so that the frequency of the label will be approximately balanced whenever the negation word exists, i.e., the data debiasing goal is satisfied.
To perform the operation $do(B=b_{k})$, we utilize the LLM itself to rewrite the data. Specifically, this causal intervention process employs the few-shot prompting method to rewrite samples within the dataset using the LLM itself. To ensure that the biased feature of the rewritten samples satisfies the intervention objectives $b_{k}$, we use the LLM itself or some metrics to check different biased features (specific details can be found in Appendix C). If the requirements are met, the rewritten data is retained. Otherwise, the rewriting process is repeated until the intervenion objectives are met or the rewriting fails to meet the objectives three times.
NLI PI SA QA
Llama3.1 MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla SFT 83.5 72.7 86.6 75.0 95.4 95.4 87.3 65.6 86.5
RAZOR 83.7 73.2 86.4 74.7 95.3 95.6 87.3 65.6 86.6
IGCIDB 83.4 75.4 86.2 77.3 95.4 95.5 87.6 66.2 86.8
Gemma2 MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla SFT 79.5 65.7 86.8 73.8 95.9 95.3 85.5 68.8 84.8
RAZOR 79.4 67.6 86.7 72.9 95.8 95.5 85.5 68.9 84.9
IGCIDB 79.1 69.0 86.5 77.6 95.8 95.5 85.6 69.1 85.2
Table 1: Comparison of IGCIDB with baselines across two LLMs on ID and transfer test sets.
Note that the number of samples used for the causal intervention process can be automatically calculated. For simplicity, we manually determine the intervention direction and the number of samples to be intervened in order to satisfy the data debiasing goal.
4 Experimental Setup
4.1 Experimental Details
In this work, we use Llama3.1-8B Dubey et al. (2024) to debias the instruction-tuning dataset, and then fine-tune Llama3.1-8B on the debiased dataset for experiments. To examine the generalizablity of the debiased dataset, Gemma2-9B Team et al. (2024) is also utilized to fine-tune upon the debiased dataset.
Our instruction-tuning dataset consists of Lima (Zhou et al., 2024) and a portion of Flan 2021 (Wei et al., 2022) datasets. Due to the time limit, we separately choose 10,000 random samples from 5 sub-datasets of the Flan 2021 dataset, including MNLI (Williams et al., 2018), QQP (Sharma et al., 2019), SST2 (Socher et al., 2013), Squadv1 (Rajpurkar, 2016), and TriviaQA (Joshi et al., 2017). These sub-datasets include four types of tasks: natural language inference (NLI), paraphrase identification (PI), sentiment analysis (SA), and question answer (QA). Since the task of the Lima dataset is multi-turn dialogue which is an open-ended task without standard answers, the biased features cannot be automatically induced by the CAL method Sun et al. (2024), and hence we apply the IGCIDB method to the other four tasks.
We utilized lora (Hu et al., 2022) for supervised fine-tuning (SFT) in our experiments. During the training process, we employed a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 1e-4 for both models. We do not use early stopping, but instead train for three epochs. More details can be seen in Appendix D.
4.2 Evaluation Tasks
We examine the effectiveness of the information gain-guided causal intervention debiasing method by investigating whether it could debias LLMs to improve the zero-shot generalizability of LLMs. Specifically, we examine the generalizability of LLMs using transfer test sets and challenge test sets.
ID Test Sets We evaluate our debiasing framework on five in-domain (ID) datasets including MNLI, QQP, SST2, Squadv1, and TriviaQA.
Transfer Test Sets We evaluate the performance of LLMs in maintaining strong transferability across datasets for the same task, using SNLI (Bowman et al., 2015), MRPC (Dolan and Brockett, 2005), IMDB (Maas et al., 2011), and Squadv2 (Rajpurkar et al., 2018) as transfer test sets for NLI, PI, SA, and QA tasks, respectively.
Challenge Test Sets We assess the robustness of models against dataset bias using challenge test sets, specifically designed by manually eliminating the dataset biases from the dataset to challenge models. With the lack of corresponding challenge datasets for QA and SA task, we use HANS McCoy et al. (2019) and PAWS Zhang et al. (2019) as challenge test sets for NLI and PI task.
Additionally, to evaluate the general ability of LLMs, we utilize the extensively used MMLU (Hendrycks et al., 2021) and BBH (Suzgun et al., 2023) datasets. For BBH datasets, we use the exact match as the evaluation metric. For other datasets, we measured the accuracy to compare the results of different methods.
4.3 Baseline Methods
We compare the information gain-guided causal intervention debiasing method with two categories of training-based methods:
Vanilla Supervised Fine-tuning Baselines We examine the performance of LLMs after vanilla supervised fine-tuning.
NLI PI
Llama3.1 MNLI HANS QQP PAWS
Vanilla SFT 83.5 71.3 86.6 63.4
RAZOR 83.7 72.5 86.4 62.3
IGCIDB 83.4 74.2 86.2 66.9
Gemma2 MNLI HANS QQP PAWS
Vanilla SFT 79.5 65.8 86.8 59.9
RAZOR 79.4 67.8 86.7 58.4
IGCIDB 79.1 70.4 86.5 62.6
Table 2: Comparison of IGCIDB with baselines across two LLMs on ID and challenge test sets.
Llama3.1 Gemma2
Methods MMLU BBH MMLU BBH
Vanilla SFT 60.2 42.6 68.8 44.9
RAZOR 59.3 41.7 68.9 44.3
IGCIDB 60.5 42.3 68.5 45.6
Table 3: Comparison of IGCIDB with baselines across two LLMs on the general ability benchmarks.
Prior Knowledge-Based Debiasing Method Because there are no training-based automatic debiasing methods for LLMs. So we choose the latest prior knowledge-based debiasing method RAZOR (Yang et al., 2025) and adapt it to the instruction tuning setting of LLMs. More details can be found in the Appendix E.
5 Experimental Results
5.1 Generalizability on Transfer Test Sets
We list the experimental results of two LLMs on ID and transfer test sets in Table 1. From which we find that:
Figure 3: The density plot curve for the model confidence of answer contradiction when experimented with Llama3.1-8B.
Figure 4: The density plot curve for the model confidence of answer contradiction when experimented with Gemma2-9B.
(1) Compared to the vanilla supervised fine-tuning method, in general, the prior knowledge-based debiasing method RAZOR shows improved performance on the transfer test sets. This indicates that by utilizing prior knowledge to debias LLM, RAZOR can improve the generalizability of LLMs. However, it can be found that the performance of RAZOR on the paraphrase identification task is decreased. This may be due to the prior knowledge that the dataset bias originates solely from specific tokens in the input is not applicable for all tasks. This demonstrates the limitations of the prior knowledge-based debiasing method and further illustrates the importance of automatic debiasing methods for LLMs.
Task Examples Gold Baseline IGCIDB
QA Scottish Parliament: … a member of the Scottish Parliament can introduce a member’s bill as a private member; or a private bill can be submitted to Parliament by an outside proposer. … Who may also submit private bills to the Parliament? An outside proposer Scottish Parliament’s member × An outside proposer ✓
QA Scottish Parliament: … a member of the Scottish Parliament can introduce a member’s bill as a private member; or a private bill can be submitted to Parliament by an outside proposer. … Who can submit member’s bill to the Parliament? Scottish Parliament’s member Scottish Parliament’s member ✓ Scottish Parliament’s member ✓
Table 4: Examples of case study. The baseline method tends to use the biased feature ‘individual’s popularity’ to answer Scottish Parliament’s member rather than an outside proposer in the first example.
(2) Compared to the vanilla supervised fine-tuning method, IGCIDB achieves consistent performance improvement on transfer test sets while maintaining performance on the in-domain datasets. This demonstrates that by leveraging the information gain-guided causal intervention, the biased features will not contribute to accurately predicting the answers for the instruction-tuning dataset. As a result, IGCIDB can effectively debias LLMs by fine-tuning on the debiased dataset to improve the generalizability of LLMs.
(3) Compared with the prior knowledge-based debiasing method RAZOR, IGCIDB shows relatively better performance on transfer test sets, which demonstrates the effectiveness of our method for debiasing LLMs. Additionally, compared to the vanilla supervised fine-tuning method, the performance of IGCIDB increases on all transfer test sets, while the performance of RAZOR decreases on the MRPC dataset. On the one hand, the performance degradation of RAZOR shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks. On the other hand, the performance improvement of IGCIDB on all transfer test sets further indicates the necessity of automatically identifying and eliminating dataset biases.
(4) In general, our method is effective for both Llama3.1-8B and Gemma2-9B. Moreover, even if fine-tuning Gemma2-9B on the dataset debiased using Llama3.1-8B, the effectiveness of debiasing to improve generalizability still remains. This demonstrates the generality of our approach in adapting to different LLMs. That is to say, once we debias a dataset using a certain LLM, we can directly use it to debias other LLMs.
5.2 Generalizability on Challenge Test Sets
To investigate the generalizability of IGCIDB on the manually debiased datasets, we examine our method on HANS and PAWS datasets. As shown in Table 2, compared to the prior knowledge-based debiasing method RAZOR and the vanilla supervised fine-tuning method, IGCIDB shows better performance on both challenge test sets while maintaining performance on the in-domain datasets, which underscores the efficacy of our method for debiasing LLMs. Moreover, compared to the vanilla supervised fine-tuning method, the performance of IGCIDB increases on both challenge test sets, while the performance of RAZOR decreases on PAWS. This shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks, and further demonstrates the necessity of automatically identifying and eliminating dataset biases.
5.3 General Ability of LLMs
Though effectiveness for debiasing LLMs, we hope that the general ability of LLMs will not be decreased compared to the vanilla supervised fine-tuning method. Therefore, we utilize MMLU and BBH datasets to examine the general ability of different methods.
As shown in Table 3, the general ability of the IGCIDB method is not decreased and even a little better than that of the vanilla supervised fine-tuning method when using Gemma2-9B for experiments. However, the performance of RAZOR is generally lower than the vanilla supervised fine-tuning method. This further indicates the effectiveness of our method.
5.4 Analysis of Model Confidence
To further investigate the effectiveness of the IGCIDB method, we use the kernel density estimation to draw the distributions of model confidence on bias-conflicting samples. Specifically, we conduct experiments on the HANS dataset, where there is a high lexical overlap rate between the premise and hypothesis across all data. McCoy et al. (2019) have demonstrated that for the NLI task, LLMs tend to predict entailment rather than contradiction when there is a high lexical overlap rate between the premise and hypothesis. Therefore, we consider the samples with label contradiction as the bias-conflicting samples.
The result is illustrated in Figure 3 and Figure 4. It is observed that the vanilla supervised fine-tuning method shows high density on the low confidence of the answer contradiction, which demonstrates that LLMs tend to use the biased feature lexical overlap for predicting entailment after vanilla supervised fine-tuning. On the contrary, our method shows a much lower density on the low confidence of the answer contradiction compared to the vanilla supervised fine-tuning method. This is because, the information gain-guided causal intervention ensures that the biased feature ‘high lexical overlap’ does not contribute to accurately predicting the answers. Consequently, LLMs will not exhibit obvious bias after fine-tuning on the debiased dataset.
5.5 Case Study
To demonstrate the efficacy of the proposed method, we present a case study for QA task in Table 4 (more case studies can be seen in the Appendix F). We compare our proposed method based on Llama3.1-8B with the baseline of the vanilla supervised fine-tuning method.
From Table 4, we can find that the baseline method tends to use the biased feature “individual’s popularity” to answer Scottish Parliament’s member in both cases. However, since the given context explicitly states “a private bill can be submitted to Parliament by an outside proposer”, the correct answer is “an outside proposer” which is less popular than “Scottish Parliament’s member” in the first example. Combining the case studies in the Appendix F, it can be concluded that our method IGCIDB, exhibits enhanced robustness compared to the baseline approach. The resilience of our method to conditions with different biased features suggests that it is well-suited for real-world environments where reliable LLMs are crucial.
6 Related Work
LLMs Debiasing Previous works demonstrate that LLMs still suffer from biases such as negation bias Mukherjee et al. (2021) and popularity bias Klimashevskaia et al. (2024). To debias LLMs, a series of methods depend on researchers’ prior knowledge to artificially detect the potential dataset biases, followed by debiasing through prompt-based regularization, instruction tuning, or preference optimization Wang et al. (2023); Ganguli et al. (2023); Tang et al. (2024). However, these methods are limited by the reliance on researchers’ prior knowledge. Additionally, due to the diversity of dataset biases Poliak et al. (2018); Schuster et al. (2019); Schick et al. (2021), it is impractical to identify them one by one manually for all tasks and datasets.
To address these issues, previous works proposed automatic debiasing methods. They train biased models to automatically extract biased features which characterizes the dataset biases Utama et al. (2020b); Du et al. (2023); Sanh et al. (2020); Lyu et al. (2023), and then some methods such as confidence regularization (Utama et al., 2020a) are utilized for regularizing the main model. However, such methods are designed for discriminative models and are difficult to adapt to generative LLMs. To automatically debias generative LLMs, Sun et al. (2024) proposed the causal-guided active learning method, which first automatically induces explainable biased features, and then debias LLM utilizing in-context learning. However, this method is based on in-context learning so that it can only suppress bias in LLMs but cannot completely eliminate it in theory. Moreover, due to the limited instruction following ability of LLMs, it can only address limited numbers of biases. For example, it can only suppress two biases when using Llama2-13b-Instruct.
In this paper, we propose an information gain-guided causal intervention debiasing framework for automatically debiasing generative LLMs. We borrow the idea from information theory and propose an information gain-guided data debiasing goal, followed by a causal intervention-based data rewriting method for debiasing instruction-tuning datasets. Based on the debiased dataset, the generalizability of LLMs can be improved using standard supervised fine-tuning.
Causal Mechanism A series of recent works have explored to debias the neural network models based on the causal mechanism. Lv et al. (2022) utilized Structural Causal Model (SCM) to model auxiliary variables and exhibit promising performance in domain generalization. Sun et al. (2024) explored the causal invariance mechanism to automatically identify biased samples and biased features. In this work, we use the causal intervention (Pearl, 2009) to automatically debias LLMs.
7 Conclusions
In this paper, we propose an information gain-guided causal intervention debiasing framework. According to the condition of unbiased dataset that the biased feature should not provide any additional information for predicting the answers, we can derive the data debiasing goal using information gain theory. Then we utilize a causal intervention-based data rewriting method for achieving the data debiasing goal. Finally, a standard supervised fine-tuning process is used to debias LLMs. Experimental results show that our approach can effectively debias LLMs to enhance their generalizability in both transfer and challenge settings.
8 Limitations
Although our method can automatically debias LLMs, the bias identification process is based on the causal-guided active learning method, which cannot deal with tasks that do not have a standard answer such as writing essays.
Appendix A Proof of the information Gain-guided Data Debiasing Goal
From Eq. 1, we can know that:
$H(Y|B)=H(Y),$ (5)
where $H$ is the information entropy. The entropy $H(Y)$ and the conditional entropy $H(Y|B)$ over all possible values of $Y$ and $B$ are defined as:
$H(Y)=\sum_{y\in Y}P(y)\log_{2}P(y),$ (6)
$H(Y|B)=\sum_{b\in B}\sum_{y\in Y}P(y,b)\log_{2}P(y|b),$ (7)
Substitute Eq. 5 and Eq. 6 into Eq. 4, we have
$\sum_{y\in Y}P(y)\log_{2}P(y)=\sum_{b\in B}\sum_{y\in Y}P(y,b)\log_{2}P(y|b),$ (8)
By using probability formulas, we have
$\sum_{b\in B}\sum_{y\in Y}P(b)P(y|b)\log_{2}P(y)=$
$\sum_{b\in B}\sum_{y\in Y}P(b)P(y|b)\log_{2}P(y|b),$ (9)
Note that $\sum_{b\in B}\sum_{y\in Y}P(b)P(y|b)$ is a public item, so we have
$\sum_{b\in B}\sum_{y\in Y}P(b)P(y|b)(\log_{2}P(y)-\log_{2}P(y|b))=0,$ (10)
Because $p(y|b)$ is always greater than or equal to $P(y)$, In order for Eq. 9 to hold, it is necessary to ensure that $P(y)=P(y|b)$ for any y and b, so that we have
$P(Y)=P(Y|B),$ (11)
and finishes the proof.
Appendix B Data Debiasing Goal Based on Different Biased Features
For each task, we choose the most general biased features identified by the CAL method for data debiasing. Specifically, for NLI and PI task, the identified biased feature is the lexical overlap between the two sentences. However, the lexical overlap rate can be any value between 0 and 1. To establish the data debiasing goal, we manually defined the value space of this biased feature: low lexical overelap (for lexical overelap rate which is less than 0.4), medium lexical overelap (between 0.4 and 0.6), low lexical overelap (bigger than 0.6). For these two tasks, the value space of the answer is easy to enumerate, so no modifications are necessary. For the SA task, the identified bias feature is whether the negation word exists in the sentence, so the data debiasing goal does not need to be adjusted as illustrated in Sec. 3.2. For the bias identification in the QA task, we choose two primary categories of questions (as shown in D): those aim at inquiring about a person and those ask for a location. From these two categories, the identified biased features are the popularity of individual and the popularity of location respectively. For the biased feature popularity of the individual, the adjustment process has been shown in Sec. 3.2. For the biased feature popularity of the location, the adjustment process is similar.
Task Examples Gold Baseline IGCIDB
SA If you’re looking for a not-so-serious mob movie, with a female as the lead, you’re in the right place. Pfieffer has acted much better than this. You can see she has matured beyond this picture. When I first picked this movie up, I expected Pfeiffer was poorly miscast, however, she plays her mob wife role to the hilt. Not a bad performance from Baldwin, either. If you don’t pay attention to the hair, you might enjoy this movie. What is the sentiment of this review? Positive Negative × Positive ✓
PI Sentence 1: However, Ann decides Peggy loses her visitation as the Bed and Breakfast is not safe. Sentence 2: Peggy decides that Ann loses her visitation because the bed and breakfast is not safe. Do these sentences mean the same thing? No Yes × No ✓
NLI Sentence 1: IDAs are special in that low-income savers receive matching funds from federal and state governments as well as private sector organizations as an incentive to save. Sentence 2: IDAs are special in that low-income savers receive differing funds from federal and state governments. Does the sentence 1 entail the sentence 2? No Yes × No ✓
Table 5: Examples of case study.
Appendix C Details of the biased feature checking process
To ensure that the biased feature of the rewritten samples satisfies the rewriting objectives, we employ different metrics to check these biased features. For lexical overlap bias, we calculate the lexical overlap rate between two input sentences. For negation word bias, we directly verify the presence of any predefined negation words in the inputs. For the popularity bias, we utilize the biased LLM to check if the individual or the location is considered popular.
Appendix D More Experimental Details
When applying the IGCIDB method to QA tasks, we focus on two primary categories of questions for data debiasing due to the CAL method’s requirement for uniformity in the question’s objective. One category involves questions aimed at inquiring about a person, while the other involves questions about a location. After data debiasing, we also integrate those unchanged data (such as the Lima dataset) into the debiased dataset. For the evaluation of the MNLI, QQP, MRPC, SST2, Squadv1, and Squadv2 datasets, we utilize the validation sets for evaluation due to the unavailability of the test sets. For the other datasets, we use the test sets for evaluation.
Appendix E Details of the RAZOR Baselines
The RAZOR method is based on the prior knowledge that the bias may originate from the specific token of the inputs. Based on this, it begins by defining a bias score to evaluate the degree of bias for each piece of data, followed by rewriting the data to achieve low bias scores for data debiasing. However, the calculation of the bias score requires to know the label space of the task, so that it cannot be applied to the QA and conversation task. As a result, when using it as a baseline, we apply this method to the MNLI, QQP, and SST2 datasets separately, and then merge them with the other unchanged instruction-tuning data to form a debiased instruction-tuning dataset for the subsequent supervised fine-tuning step.
Appendix F More Case Studies
Here are case studies for the SA, PI, and NLI tasks. As illustrated in Table. 5, the first case contains negation words in the review, yet the sentiment is positive. The IGCIDB method accurately predict its sentiment. However, the baseline method relies on the biased feature negation word to predict negative sentiment. This highlights the effectiveness of our method. In the PI and NLI task examples, there is a high lexical overlap rate between the two sentences, leading the baseline method to make incorrect predictions based on this biased feature. However, our method accurately predicts the correct answers for both samples. Combining the case studies in Sec. 5.5, it can be concluded that IGCIDB exhibits improved robustness compared to the baseline approach.